The sum of slopes the line x²-2cxy -7y²= 0 is 4times their product c has the value

Grade:12th pass

1 Answers

Arun
6 years ago
Let the lines are
L1:
y = m1x
And L2:
y = m2x
Then m1 + m2 = -2c/7
m1* m2= -1/7
Given
m1 +m2 = 4 m1*m2
Put the values and you will get answer
as c = 2